The length of a rectangle is 3 times its width. If the perimeter is 64 yards what is the area?
Answer by checkley77(12844)

L=3W
P=2L+2W
64=2*3W+2W
64=6W+2W
64=8W
8W=64
W=64/8
W=8 ANS. FOR THE WIDTH.
L=3*8=24 ANS. FOR THE LENGTH.
PROOF:
64=2*24+2*8
64=48+16
64=64